# Joint Optimization on Uplink OFDMA and MU-MIMO for IEEE 802.11ax: Deep Hierarchical Reinforcement Learning Approach

### v1

@@ Abstract
This letter tackles a joint user scheduling, frequency resource allocation (USRA), multi-input-multi-output mode selection (MIMO MS) between single-user MIMO and multi-user (MU) MIMO, and MU-MIMO user selection problem, integrating uplink orthogonal frequency division multiple access (OFDMA) in IEEE 802.11ax. Specifically, we focus on unsaturated traffic conditions where users’ data demands fluctuate. In unsaturated traffic conditions, considering packet volumes per user introduces a combinatorial problem, requiring the simultaneous optimization of MU-MIMO user selection and RA along the time-frequency-space axis. Consequently, dealing with the combinatorial nature of this problem, characterized by a large cardinality of unknown variables, poses a challenge that conventional optimization methods find nearly impossible to address. In response, this letter proposes an approach with deep hierarchical reinforcement learning (DHRL) to solve the joint problem. Rather than simply adopting off-the-shelf DHRL, we tailor the DHRL to the joint USRA and MS problem, thereby significantly improving the convergence speed and throughput. Extensive simulation results show that the proposed algorithm achieves significantly improved throughput compared to the existing schemes under various unsaturated traffic conditions.
@@ I Introduction
While applying off-the-shelf DHRL directly makes the learning process nearly impractical due to agents’ explosive action space size, we refine sub-agents to align with the standard-defined frequency band utilization, markedly improving the practicality and performance of the system. In addition, we propose a state update process based on channels’ semiorthogonality, which leads to a huge performance gain. Simulation results demonstrate that the proposed DHRL algorithm achieves significantly higher throughput under unsaturated traffic conditions than the existing schemes.
@@ III-B Structure and parameters of the proposed DHRL
  • Goal: The goal at the time step $t$ is defined by $g_{t}\in\mathcal{G}$, where $\mathcal{G}$ is the goal space including all possible cases of RA according to the IEEE 802.11ax standard [1].
@@ III-C Sub-agent design for MU-MIMO
To reduce the action space sizes of sub-agents to a level comparable to SU-MIMO, we define the decision round $\tilde{t}$ within a time step $t$ and design sub-agents that repeatedly select one STA at each decision round for at most $\lfloor N_{\text{R}}/N_{\text{T}}\rfloor$ times. Additionally, at each decision round, the agent can take a ”break” option, which allows for the early stop of the action selection procedure to select fewer actions than $\lfloor N_{\text{R}}/N_{\text{T}}\rfloor$. Then, we significantly reduce the action space sizes from $\sum_{i=1}^{\lfloor N_{\text{R}}/N_{\text{T}}\rfloor}$ (${K}\choose{i}$ to $K+1$, which boosts the convergence speed of the sub-agents and the improvement of network performance.
We redefine an action as $\tilde{a}_{\tilde{t}}^{(l,i)}\in\mathcal{K}\cup\{\emptyset\}=\tilde{\mathcal{A}}$, where $\emptyset$ represents the break sign, and $\tilde{\mathcal{A}}$ is the newly defined action space. Across multiple decision rounds, each sub-agent iteratively selects actions. Subsequently, the selected actions are aggregated to form $a_{t}^{(l,i)}=\{\tilde{a}_{0}^{(l,i)},\ldots,\tilde{a}_{T-1}^{(l,i)}\}$, where $T$ is the terminal time step. After decision rounds finish, each sub-agent constitutes $\mathbf{X}_{t}$ with $a_{t}^{(l,i)}$ executes $\mathbf{X}_{t}$, and obtains intrinsic reward, $r_{t}^{(l,i)}=\sum_{k\in\mathcal{K}}R_{l,i}^{(k)}(\mathbf{X}_{t},\bm{\Gamma}_{t},\mathbf{p}_{t})$.
@@ III-D1 Initialization
We introduce the Deep Q-network (DQN) as the learning framework of the master agent and sub-agents. Thus, we utilize a parameter $\theta_{\text{M}}$ that defines an action-value function $Q(\mathcal{S},g;\theta_{\text{M}})$ for the master agent and parameters $\theta_{\text{S}}^{(l)}$ that define action-value functions $Q(s,a;\theta_{\text{S}}^{(l)},g)$ for the $l$-th level sub-agent, where $l$ ranges from 00 to $L$. In addition, we initialize replay memories $\mathcal{D}_{\text{M}}$ for the master agent to capacity $C_{\text{M}}$ and $\mathcal{D}_{\text{S}}^{(l)}$ for the $l$-th sub-agent to capacity $C_{\text{S}}^{(l)}$.
@@ III-D2 Experience collection
The master agent first selects a goal $g_{t}$ with a $\epsilon$-greedy policy, where $\epsilon$ decays from 1.0 to 0.1. For $[\text{RU}(l_{0},i_{0})$, $\text{RU}(l_{1},i_{1}),\ldots,\text{RU}(l_{d},i_{d})]$ determined by the selected goal $g_{t}$, each sub-agent selects an action $a_{t}^{(l_{j},i_{j})}$. Algorithm 2 shows the action selection process in the sub-agents. If a sub-agent supports MU-MIMO mode, the sub-agent repeatedly appends an action $\tilde{a}_{\tilde{t}}^{(l_{j},i_{j})}$ to $a^{(l_{j},i_{j})}$ for at most $\lfloor N_{\text{R}}/N_{\text{T}}\rfloor$ decision rounds. At each decision round $\tilde{t}$, the sub-agent evaluates the reward difference $r_{\tilde{t}}^{(l_{j},i_{j})}-r_{\tilde{t}-1}^{(l_{j},i_{j})}$ and stores transition $\left(\tilde{s}_{\tilde{t}}^{(l_{j},i_{j})},\tilde{a}_{\tilde{t}}^{(l_{j},i_{j})},r_{\tilde{t}}^{(l_{j},i_{j})}-r_{\tilde{t}-1}^{(l_{j},i_{j})},\tilde{s}_{\tilde{t}+1}^{(l_{j},i_{j})}\right)$ in $\mathcal{D}_{\text{S}}^{(l_{j},i_{j})}$. The master agent also stores transition $\left(\mathcal{S}_{t},g_{t},\hat{r}_{t},\mathcal{S}_{t+1}\right)$ in $\mathcal{D}_{\text{M}}$ at each time-step $t$.
@@ III-D3 Updating model parameters
The proposed deep hierarchical Q-network (DHQN)-based USRA algorithm is described in Algorithm 1.
@@
5:   if $\tilde{a}_{\tilde{t}}^{(l,i)}=\emptyset$ then break  end if
@@ III-E Complexity analysis
We find that the integrated part from the branch network for CSI to the fusion network dominates the majority of computational complexity. Thus, we here focus on the computational complexity of such a part. We first define the network depth of the integrated part as $d$. With the input size of $K\cdot 2^{L/2}$, the computational complexity of the integrated part is $\mathcal{O}(dK^{2}2^{L})$. In addition, the state update process in (4) requires $\mathcal{O}(MK)$ computation, where $M$ is the maximum iteration level. Consequently, the computational complexity of the proposed DHRL algorithm reads $\mathcal{O}(dMK^{3}2^{L})$. On the other hand, the conventional work [2] requires $\mathcal{O}(MK^{2}4^{L})$. While our DHRL algorithm requires $dK$ times more computational effort, it benefits exponentially from wider bandwidths, as reflected by the more favorable exponential term $2^{L}$ compared to $4^{L}$ in the conventional work.
@@ IV Simulation Results
  • SINR-based scheduling (fixed RA): It maximizes the throughput of STAs by optimizing US but with fixed RA. If MU-MIMO and OFDMA are utilized, the algorithm chooses level $l=\min(L-2,\lfloor\log_{2}(K\cdot N_{\text{R}}/N_{\text{T}})\rfloor)$ [2].
@@
In Fig. 5, we measure the throughput of the proposed and baseline methods by varying the number of STAs and the packet arrival rate. The SINR-based scheduling algorithm demonstrates superior throughput compared to the SINR-based scheduling and buffer-based scheduling methods with fixed RA in addressing the USRA problem, incorporating considerations for MU-MIMO and MIMO MS to find near-optimal solutions under saturated traffic conditions. However, a limitation of the SINR-based scheduling method is its lack of integration with the buffer status of all STAs. In contrast, our method surpasses baseline methods by incorporating not only MU-MIMO and MIMO MS but also the buffer status of all STAs, leading to enhanced throughput.
@@
[…]

### v2

@@ Abstract
This letter tackles a joint user scheduling, frequency resource allocation (USRA), multi-input-multi-output mode selection (MIMO MS) between single-user MIMO and multi-user (MU) MIMO, and MU-MIMO user selection problem, integrating uplink orthogonal frequency division multiple access (OFDMA) in IEEE 802.11ax. Specifically, we focus on unsaturated traffic conditions where users’ data demands fluctuate. In unsaturated traffic conditions, considering packet volumes per user introduces a combinatorial problem, requiring the simultaneous optimization of MU-MIMO user selection and RA along the time-frequency-space axis. Consequently, dealing with the combinatorial nature of this problem, characterized by a large cardinality of unknown variables, poses a challenge that conventional optimization methods find nearly impossible to address. In response, this letter proposes an approach with deep hierarchical reinforcement learning (DHRL) to solve the joint problem. Rather than simply adopting off-the-shelf DHRL, we tailor the DHRL to the joint USRA and MS problem, thereby significantly improving the convergence speed and throughput.
@@ I Introduction
While applying off-the-shelf DHRL directly makes the learning process nearly impractical due to agents’ explosive action space size, we refine sub-agents to align with the standard-defined frequency band utilization, markedly improving the practicality and performance of the system. In addition, we propose a state update process based on channels’ semiorthogonality, which leads to a huge performance gain.
@@ III-B Structure and parameters of the proposed DHRL
Figure 5: RU combination table corresponding to goals in a 20 MHz bandwidth channel.
@@
[…]
@@ III-C Sub-agent design for MU-MIMO
The huge size of previously defined action space stems from the practice of contemplating every possible permutation for MU-MIMO user selection at once. Instead, we propose a sequential user selection approach where $\lfloor N_{\text{R}}/N_{\text{T}}\rfloor$ decision rounds are sequentially conducted within a single time step, during which sub-agents select one STA per round. To account for scenarios where selecting fewer users than $\lfloor N_{\text{R}}/N_{\text{T}}\rfloor$ is preferable, sub-agents can choose a ‘break’ option in any decision round, allowing for the early termination of the selection process. Via the proposed user selection approach, we significantly reduce the action space sizes from $\sum_{i=1}^{\lfloor N_{\text{R}}/N_{\text{T}}\rfloor}$ (${K}\choose{i}$ to $K+1$, which boosts the convergence speed of sub-agent training and improves network performance.
We redefine an action as $\tilde{a}_{\tilde{t}}^{(l,i)}\in\mathcal{K}\cup\{break\}=\tilde{\mathcal{A}}$, where $break$ represents the break option, and $\tilde{\mathcal{A}}$ is the newly defined action space. Across multiple decision rounds, each sub-agent iteratively selects actions. Subsequently, the selected actions are aggregated to form $a_{t}^{(l,i)}=\{\tilde{a}_{0}^{(l,i)},\ldots,\tilde{a}_{T-1}^{(l,i)}\}$, where $T$ is the terminal time step. After decision rounds finish, each sub-agent constitutes $\mathbf{X}_{t}$ with $a_{t}^{(l,i)}$ executes $\mathbf{X}_{t}$, and obtains intrinsic reward, $r_{t}^{(l,i)}=\sum_{k\in\mathcal{K}}R_{l,i}^{(k)}(\mathbf{X}_{t},\bm{\Gamma}_{t},\mathbf{p}_{t})$.
@@ III-D DHRL Training Process
We introduce the Deep Q-network (DQN) as the learning framework of the master and sub-agents. Algorithm 1 describes the proposed deep hierarchical Q-network (DHQN)-based USRA. With the newly-defined action $\tilde{a}$ and action space $\tilde{\mathcal{A}}$ in section III-C, we train a DHRL model as follows.
@@ III-D1 Initialization
We utilize a parameter $\theta_{\text{M}}$ that defines an action-value function $Q(\mathcal{S},g;\theta_{\text{M}})$ for the master agent and parameters $\theta_{\text{S}}^{(l)}$ that define action-value functions $Q(s,a;\theta_{\text{S}}^{(l)},g)$ for the $l$-th level sub-agent, where $l$ ranges from 00 to $L$. In addition, we initialize replay memories $\mathcal{D}_{\text{M}}$ for the master agent to capacity $C_{\text{M}}$ and $\mathcal{D}_{\text{S}}^{(l)}$ for the $l$-th sub-agent to capacity $C_{\text{S}}^{(l)}$.
@@ III-D2 Experience collection
The master agent first selects a goal $g_{t}$ with a $\epsilon$-greedy policy, where $\epsilon$ decays from 1.0 to 0.1. For $[\text{RU}(l_{0},i_{0})$, $\text{RU}(l_{1},i_{1}),\ldots,\text{RU}(l_{d},i_{d})]$ determined by the selected goal $g_{t}$, each sub-agent selects an action $a_{t}^{(l_{j},i_{j})}$. Algorithm 2 shows the action selection process in the sub-agents. A sub-agent supporting MU-MIMO repeatedly appends an action $\tilde{a}_{\tilde{t}}^{(l_{j},i_{j})}$ to $a^{(l_{j},i_{j})}$ for at most $\lfloor N_{\text{R}}/N_{\text{T}}\rfloor$ decision rounds. At each decision round $\tilde{t}$, the sub-agent evaluates the reward difference $r_{\tilde{t}}^{(l_{j},i_{j})}-r_{\tilde{t}-1}^{(l_{j},i_{j})}$ and stores transition $\left(\tilde{s}_{\tilde{t}}^{(l_{j},i_{j})},\tilde{a}_{\tilde{t}}^{(l_{j},i_{j})},r_{\tilde{t}}^{(l_{j},i_{j})}-r_{\tilde{t}-1}^{(l_{j},i_{j})},\tilde{s}_{\tilde{t}+1}^{(l_{j},i_{j})}\right)$ in $\mathcal{D}_{\text{S}}^{(l_{j},i_{j})}$. The master agent also stores transition $\left(\mathcal{S}_{t},g_{t},\hat{r}_{t},\mathcal{S}_{t+1}\right)$ in $\mathcal{D}_{\text{M}}$ at each time-step $t$.
@@ III-D3 Updating model parameters
5:   if $\tilde{a}_{\tilde{t}}^{(l,i)}=break$ then break  end if
@@ III-E Complexity analysis
We find that the integrated part from the branch network for CSI to the fusion network dominates the majority of computational complexity. Thus, we focus on the computational complexity of such a part. Defining the network depth of the integrated part as $d$, with the input size of $K\cdot 2^{L/2}$, the computational complexity is $\mathcal{O}(dK^{2}2^{L})$. In addition, the state update process in (4) requires $\mathcal{O}(MK)$ computation, where $M$ is the maximum iteration level. Then, the computational complexity of the proposed DHRL algorithm reads $\mathcal{O}(dMK^{3}2^{L})$. On the other hand, the conventional work [2] requires $\mathcal{O}(MK^{2}4^{L})$. While our DHRL algorithm requires $dK$ times more computational effort, it benefits exponentially from wider bandwidths, as reflected by the more favorable exponential term $2^{L}$ compared to $4^{L}$ in the conventional work.
@@ IV Simulation Results
  • SINR-based scheduling (fixed RA): It maximizes the STAs’ throughput by optimizing US but with fixed RA. If MU-MIMO and OFDMA are utilized, the algorithm chooses level $l=\min(L-2,\lfloor\log_{2}(K\cdot N_{\text{R}}/N_{\text{T}})\rfloor)$ [2].
@@
Figure 8: Throughput across various numbers of antennas for AP and STAs. (a) $N_{\text{T}}=1$. (2) $N_{\text{T}}=2$.
In Fig. 6, we measure the throughput by varying the number of STAs and packet arrival rates. The SINR-based scheduling algorithm shows higher throughput than both SINR and buffer-based methods with fixed RA in addressing the USRA problem, especially under saturated traffic conditions with MU-MIMO and MIMO MS. However, it does not account for the buffer status of STAs. Our method, which includes MU-MIMO, MIMO MS, and STA buffer statuses, outperforms the baselines. At an extreme frame rate of 10000 frames/s, our method achieves throughput levels comparable to or better than the baselines, even in saturated conditions, as shown in Fig. 7. In Fig. 8, we assess the throughput as the number of antennas on the AP and STAs increases. Our method consistently surpasses the baseline in all antenna configurations.